Find f(x) if f(x) = 177 and f(1) = 1 Evaluate the definite integral -, xe** dx
Answer #1

Ans: f(x) = Inca) Sf(x) = sin(x) dx. f(0) = In(x) da- x Let u= n(x) du= the dx fla (4.dy) f(a). +C ff86) [10(x)] 2 + C 2 Atxex² dx Solving indefinite integral first) że ex² dx- X Let u=x2 du = 2x.dx => du · xp to > e. du 2 7 Seudu.. => aza 2 So, &